# Cellular, Cell-less, and Everything in Between: A Unified Framework for Utility Region Analysis in Wireless Networks

### v2

@@ I-C Relation to existing studies
Convexity of the SINR region for systems with affine interference functions has been studied in [13, Ch. 5][24]. In those studies, convexity is only established if the interference matrices described later in Sect. III are symmetric, which is an assumption unlikely to hold in practice. Our analysis lifts the symmetry assumption and incorporates power constraints. Furthermore, many of our results extend to nonlinear systems, which is essential for modeling networks employing optimal beamforming strategies.
@@
The results that follow are also related to those in [9, 10, 26, 11, 27], which are studies focusing on the sum-rate maximization problem using interference models that do not take into account self-interference or beamforming gain uncertainty, which has become a critical aspect in recent studies on extremely large MIMO arrays and distributed MIMO (and, in particular, cell-less) systems. Instead, our focus is on establishing sufficient conditions for the convexity of the entire SINR and rate regions of these modern interference models. By doing so, we identify a particular class of problems where the sum-rate maximization can be posed as a convex optimization problem. Notably, we do not impose symmetry on the interference patterns [12, Sect. V.A] or assume networks operating in the interference-free or in the low signal-to-noise ratio (SNR) regime [9, Sect. IV.B]. Our sufficient conditions for convexity of the SINR region (which are related to the concept of inverse Z-matrices, as also uncovered in [27] for interference models in digital subscriber lines) can identify convexity in scenarios where interference in some links is strong. Furthermore, unlike [9, 10, 26, 11, 27], we can cope with power constraints that are not necessarily convex polytopes, and we use the spectral radius of nonlinear mappings to study interference of modern network models. This approach enables the presentation of the core ideas with compact and straightforward notation, significantly enhancing clarity and accessibility after the technical foundations have been established, even if we consider complex beamforming strategies. See, for example, Remark 3 in Sect. III-B.
@@ II-A Notation and standard definitions
A diagonal matrix with positive components in the diagonal is a positive diagonal matrix, and the set of nonnegative and positive diagonal matrices of dimension $N\times N$ are denoted by, respectively, $\mathcal{D}_{+}^{N\times N}$ and $\mathcal{D}_{++}^{N\times N}$. By $\mathrm{diag}(x_{1},\ldots,x_{N})=\mathrm{diag}({\bm{x}})$, where ${\bm{x}}=(x_{1},\ldots,x_{N})\in{\mathbb{R}}^{N}$, we denote a ${\mathbb{R}}^{N\times N}$ diagonal matrix having the scalars $x_{1},\ldots,x_{N}$ in the diagonal entries. We write the transpose of a matrix $\bm{M}$} as ${\bm{M}}^{t}$, and the matrix exponential as $e^{\bm{M}}$. Any matrix ${\bm{M}}\in{\mathbb{R}}^{N\times N}$ can be interpreted as the linear mapping ${\mathbb{R}}^{N}\to{\mathbb{R}}^{N}:{\bm{x}}\mapsto{\bm{Mx}}$ (standard basis), so we write ${\bm{M}}T$ to denote the mapping $\mathcal{X}\to{\mathbb{R}}^{N}:{\bm{x}}\mapsto({\bm{M}}\circ T)({\bm{x}})={\bm{M}}(T({\bm{x}}))$, where $T:\mathcal{X}\to\mathcal{X}$ and $\mathcal{X}\subset{\mathbb{R}}^{N}$. A matrix ${\bm{M}}\in{\mathbb{R}}^{N\times N}$ is called a Z-matrix if it has nonpositive off-diagonal components [29][30]. If a matrix ${\bm{M}}\in{\mathbb{R}}^{N\times N}$ has an inverse ${\bm{M}}^{-1}$ that it is a Z-matrix, then $\bm{M}$} is said to be an inverse Z-matrix. A matrix ${\bm{M}}\in{\mathbb{R}}^{N\times N}$ is called an M-matrix if it can be expressed as ${\bm{M}}=\alpha{\bm{I}}-{\bm{P}}$ for some ${\bm{P}}\in{\mathbb{R}}^{N\times N}_{+}$ and $\alpha>\rho$, where $\rho$ is the spectral radius of $\bm{P}$}, in the conventional sense in linear algebra. If ${\bm{M}}^{-1}$ is an M-matrix, then $\bm{M}$} is called an inverse M-matrix.
@@
For convenience, given $(N,M)\in{\mathbb{N}}\times{\mathbb{N}}$ and two sets $\mathcal{X}\subset{\mathbb{R}}^{N}$ and $\mathcal{Y}\subset{\mathbb{R}}^{M}$, we say that the mapping $f:\mathcal{X}\to\mathcal{Y}$ is a function only if $M=1$ (NOTE: if $M=1$, $f$ is both a function and a mapping). A mapping $f:\mathcal{X}\subset{\mathbb{R}}^{N}\to\mathcal{Y}\subset{\mathbb{R}}^{M}$ is continuous if, for every sequence $({\bm{x}}_{n})_{n\in{\mathbb{N}}}$ in $\mathcal{X}$ converging to ${\bm{x}}^{\star}\in\mathcal{X}$, the sequence $(f({\bm{x}}_{n}))_{n\in{\mathbb{N}}}$ in $\mathcal{Y}$ converges to $f({\bm{x}}^{\star})\in\mathcal{Y}$. Let ${S}\subset{\mathbb{R}}^{N}$ be a nonempty convex set. A function $f:{S}\to{\mathbb{R}}$ is convex if $(\forall\alpha\in~]0,1[)(\forall{\bm{x}}\in{S})(\forall{\bm{y}}\in{S})~f(\alpha{\bm{x}}+(1-\alpha){\bm{y}})\leq\alpha f({\bm{x}})+(1-\alpha)f({\bm{y}}),$ and recall that its level set $\{{\bm{x}}\in{S}~|~f({\bm{x}})\leq\beta\}$ is a (possibly empty) convex set for every $\beta\in{\mathbb{R}}$. A function $f:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$ is monotonic or monotone if $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})(\forall{\bm{y}}\in{\mathbb{R}}_{+}^{N})~{\bm{x}}\geq{\bm{y}}\Rightarrow f({\bm{x}})\geq f({\bm{y}})$; scalable if $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})(\forall\alpha>1)~f(\alpha{\bm{x}})<\alpha f({\bm{x}})$; and positively homogeneous if $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})(\forall\alpha>0)~f(\alpha{\bm{x}})=\alpha f({\bm{x}})$. Following standard terminology in the wireless literature, we introduce below two classes of mappings that have played a major role in network optimization.
@@
In the above definition, we assume everywhere continuity to avoid technical digressions. Furthermore, we have lifted some of original restrictions related to the domain and codomain of general interference functions. For example, unlike the original definition in [17, 18], the function ${\mathbb{R}}_{+}\to{\mathbb{R}}_{+}:x\mapsto 0$ is a general interference function according to our definition. It is well known that monotonicity and scalability imply positivity of a mapping in the entire nonnegative cone ${\mathbb{R}}_{+}^{N}$, so we often use ${\mathbb{R}}_{++}^{N}$ to denote the codomain of a standard interference mapping to emphasize this fact. From an application perspective, an important proper subclass of standard interference mappings is the class of positive concave mappings (see [31, Proposition 1]); i.e., mappings $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}$ such that $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})(\forall{\bm{y}}\in{\mathbb{R}}_{+}^{N})(\forall\alpha\in~]0,1[)~T(\alpha{\bm{x}}+(1-\alpha){\bm{y}})\geq\alpha T({\bm{x}})+(1-\alpha)T({\bm{y}}).$
The set of fixed points of a mapping $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ is denoted by $\mathrm{Fix}(T):=\{{\bm{x}}\in{\mathbb{R}}^{N}_{+}\mid T({\bm{x}})={\bm{x}}\}$, and the mapping defined below plays a crucial role in the analysis of the fixed point set of standard interference mappings [23, Proposition 4] (Fact 9 in the Supplemental Material):
@@
([29, Definition 3.2] Spectral radius) Let $G:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}^{N}_{+}$ be a general interference mapping. The spectral radius of $G$ is defined to be
@@
We recall that there exist simple numerical schemes able to compute the spectral radius [23, Remark 3][34]. In particular, if the general interference mapping is linear, then it can be written as ${\bm{x}}\mapsto{\bm{Mx}}$ for some nonnegative matrix ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$, and we can use classical results from Perron-Frobenius theory to confirm that $\rho({\bm{M}})$ in Definition 3, with $\bm{M}$} interpreted as a linear mapping, coincides with the usual notion of the spectral radius of the matrix $\bm{M}$} in linear algebra.
Having established the key terminology and concepts, we can proceed with the main contributions of this study.
@@
In many resource allocation problems, the question of whether given utilities (as described in Sect. I-A) can be satisfied with power constraints involving monotone norms can be answered by computing conditional eigenvalues of standard interference mappings [23, Proposition 5]. However, handling conditional eigenvalues can be cumbersome because we need to keep track of both the mapping and the norm. To avoid this issue, we begin by linking the conditional eigenvalue of standard interference mappings to the spectral radius of corresponding general interference mappings (Proposition 1). We then proceed to establish many connections between the spectral radius of these general interference mappings and monotone norms, culminating in a key result (Proposition 3) from which convexity of the SINR region in wireless networks emerges as a corollary. This result allows us to leverage a substantial body of linear algebra literature to gain deeper insights into interference patterns in wireless systems, as detailed in the next sections. Before deriving these results, we first need to impose additional structure on the standard interference mappings under consideration.
A well-known property that is widely used in the wireless literature is that taking the pointwise minimum of a finite family of standard interference functions results in a standard interference function [22]. However, this property does not necessarily hold for the pointwise infimum of an infinite family of standard interference functions. Nevertheless, in wireless systems, the families have additional structure. For example, they are typically uniformly bounded away from zero because of the presence of noise (see [19, Eq. (32)]), and in such cases we can prove that the pointwise infimum retains the properties of standard interference functions, as demonstrated in the next lemma for a particular family of standard interference functions (concrete examples in the wireless domain are shown later in Sect. III).
@@
Let $\mathcal{Y}$ be a nonempty set. Suppose that, for each $y\in\mathcal{Y}$, $h_{y}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$ is a general interference function and ${u}_{y}\in{\mathbb{R}}_{++}$ a positive scalar. Assume that there exists $\delta>0$ satisfying $(\forall y\in\mathcal{Y})~{u}_{y}\geq{\delta}>{0}$; i.e., the set $\{u_{y}\in{\mathbb{R}}_{++}~|~y\in\mathcal{Y}\}\subset{\mathbb{R}}_{++}$ is bounded away from zero. Define a function $t:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}$ by
@@
Then $t$ in (6) is a standard interference function, and $t_{\|\cdot\|}$ in (7) is a general interference function.
@@
Owing to the space limitation, we refer readers to the Supplementary Material for the proof. ∎
We use the functions in (6) and (7) in Lemma 1 to define the following proper subclass of standard interference mappings:

### v3

@@ I-C Relation to existing studies
Convexity of the SINR region for systems with affine interference functions has been studied in [35, Ch. 5][36]. In those studies, convexity is only established if the interference matrices described later in Sect. III are symmetric, which is an assumption unlikely to hold in practice. Our analysis lifts the symmetry assumption and incorporates power constraints. Furthermore, many of our results extend to nonlinear systems, which is essential for modeling networks employing optimal beamforming strategies.
@@
The results that follow are also related to those in [37, 38, 16, 44, 40], which are studies focusing on the sum-rate maximization problem using interference models that do not take into account self-interference or beamforming gain uncertainty, which has become a critical aspect in recent studies on extremely large MIMO arrays and distributed MIMO (and, in particular, cell-less) systems. Instead, our focus is on establishing sufficient conditions for the convexity of the entire SINR and rate regions of these modern interference models. By doing so, we identify a particular class of problems where the sum-rate maximization can be posed as a convex optimization problem. Notably, we do not impose symmetry on the interference patterns [14, Sect. V.A] or assume networks operating in the interference-free or in the low signal-to-noise ratio (SNR) regime [37, Sect. IV.B]. Our sufficient conditions for convexity of the SINR region (which are related to the concept of inverse Z-matrices, as also uncovered in [40] for interference models in digital subscriber lines) can identify convexity in scenarios where interference in some links is strong. Furthermore, unlike [37, 38, 16, 44, 40], we can cope with power constraints that are not necessarily convex polytopes, and we use the spectral radius of nonlinear mappings to study interference of modern network models. This approach enables the presentation of the core ideas with compact and straightforward notation, significantly enhancing clarity and accessibility after the technical foundations have been established, even if we consider complex beamforming strategies. See, for example, Remark 1 in Sect. II-B and Remark 5 in Sect. III-C.
@@ II-A Notation and standard definitions
A diagonal matrix with positive components on the diagonal is a positive diagonal matrix, and the set of nonnegative and positive diagonal matrices of dimension $N\times N$ are denoted by, respectively, $\mathcal{D}_{+}^{N\times N}$ and $\mathcal{D}_{++}^{N\times N}$. By $\mathrm{diag}(x_{1},\ldots,x_{N})=\mathrm{diag}({\bm{x}})$, where ${\bm{x}}=(x_{1},\ldots,x_{N})\in{\mathbb{R}}^{N}$, we denote an ${\mathbb{R}}^{N\times N}$ diagonal matrix having the scalars $x_{1},\ldots,x_{N}$ on the diagonal entries. We write the transpose of a matrix $\bm{M}$} as ${\bm{M}}^{t}$, and the matrix exponential as $e^{\bm{M}}$. Any matrix ${\bm{M}}\in{\mathbb{R}}^{N\times N}$ can be interpreted as a linear mapping ${\mathbb{R}}^{N}\to{\mathbb{R}}^{N}:{\bm{x}}\mapsto{\bm{Mx}}$ (standard basis), so we write ${\bm{M}}T$ to denote the mapping $\mathcal{X}\to{\mathbb{R}}^{N}:{\bm{x}}\mapsto({\bm{M}}\circ T)({\bm{x}})={\bm{M}}(T({\bm{x}}))$, where $T:\mathcal{X}\to\mathcal{X}$ and $\mathcal{X}\subset{\mathbb{R}}^{N}$. A matrix ${\bm{M}}\in{\mathbb{R}}^{N\times N}$ is called a Z-matrix if it has nonpositive off-diagonal components [28][19]. If a matrix ${\bm{M}}\in{\mathbb{R}}^{N\times N}$ has an inverse ${\bm{M}}^{-1}$ that is a Z-matrix, then $\bm{M}$} is said to be an inverse Z-matrix. A matrix ${\bm{M}}\in{\mathbb{R}}^{N\times N}$ is called an M-matrix if it can be expressed as ${\bm{M}}=\alpha{\bm{I}}-{\bm{P}}$ for some ${\bm{P}}\in{\mathbb{R}}^{N\times N}_{+}$ and $\alpha>\rho$, where $\rho$ is the spectral radius of $\bm{P}$}, in the conventional sense in linear algebra. If ${\bm{M}}^{-1}$ is an M-matrix, then $\bm{M}$} is called an inverse M-matrix.
@@
For convenience, given $(N,M)\in{\mathbb{N}}\times{\mathbb{N}}$ and two sets $\mathcal{X}\subset{\mathbb{R}}^{N}$ and $\mathcal{Y}\subset{\mathbb{R}}^{M}$, we say that the mapping $f:\mathcal{X}\to\mathcal{Y}$ is a function only if $M=1$ (NOTE: if $M=1$, $f$ is both a function and a mapping). A mapping $f:\mathcal{X}\subset{\mathbb{R}}^{N}\to\mathcal{Y}\subset{\mathbb{R}}^{M}$ is continuous if, for every sequence $({\bm{x}}_{n})_{n\in{\mathbb{N}}}$ in $\mathcal{X}$ converging to ${\bm{x}}^{\star}\in\mathcal{X}$, the sequence $(f({\bm{x}}_{n}))_{n\in{\mathbb{N}}}$ in $\mathcal{Y}$ converges to $f({\bm{x}}^{\star})\in\mathcal{Y}$. Let ${S}\subset{\mathbb{R}}^{N}$ be a nonempty convex set. A function $f:{S}\to{\mathbb{R}}$ is convex if $(\forall\alpha\in~]0,1[)(\forall{\bm{x}}\in{S})(\forall{\bm{y}}\in{S})~f(\alpha{\bm{x}}+(1-\alpha){\bm{y}})\leq\alpha f({\bm{x}})+(1-\alpha)f({\bm{y}}),$ and recall that its level set $\{{\bm{x}}\in{S}~|~f({\bm{x}})\leq\beta\}$ is a (possibly empty) convex set for every $\beta\in{\mathbb{R}}$. A function $f:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$ is monotonic or monotone if $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})(\forall{\bm{y}}\in{\mathbb{R}}_{+}^{N})~{\bm{x}}\leq{\bm{y}}\Rightarrow f({\bm{x}})\leq f({\bm{y}})$; scalable if $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})(\forall\alpha>1)~f(\alpha{\bm{x}})<\alpha f({\bm{x}})$; and positively homogeneous if $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})(\forall\alpha>0)~f(\alpha{\bm{x}})=\alpha f({\bm{x}})$. Following standard terminology in the wireless literature, we introduce below two classes of mappings that have played a major role in network optimization.
@@
We recall from [7] that monotonicity and either scalability or homogeneity on the cone ${\mathbb{R}}_{+}^{N}$ implies continuity on ${\mathbb{R}}_{++}^{N}$, but not necessarily on the boundary of the domain ${\mathbb{R}}_{+}^{N}$. For this reason, we assume continuity everywhere in the above definition to avoid technical digressions. Furthermore, we have lifted some of original restrictions related to the domain and codomain of general interference functions. For example, unlike the original definition in [4, 33], the function $f:{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}:x\mapsto 0$ is a general interference function according to our definition. It is well known that monotonicity and scalability imply positivity of a mapping in the entire nonnegative cone ${\mathbb{R}}_{+}^{N}$, so we often use ${\mathbb{R}}_{++}^{N}$ to denote the codomain of a standard interference mapping to emphasize this fact. From an application perspective, an important proper subclass of standard interference mappings is the class of continuous positive concave mappings (see [11, Proposition 1]); i.e., continuous mappings $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}$ such that $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})(\forall{\bm{y}}\in{\mathbb{R}}_{+}^{N})(\forall\alpha\in~]0,1[)~T(\alpha{\bm{x}}+(1-\alpha){\bm{y}})\geq\alpha T({\bm{x}})+(1-\alpha)T({\bm{y}}).$
The set of fixed points of a mapping $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ is denoted by $\mathrm{Fix}(T):=\{{\bm{x}}\in{\mathbb{R}}^{N}_{+}\mid T({\bm{x}})={\bm{x}}\}$, and the mapping defined below plays a crucial role in the analysis of the fixed point set of standard interference mappings (see, for example, Fact 9 in the Supplemental Material):
@@
([28, Definition 3.2] Spectral radius) Let $G:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}^{N}_{+}$ be a general interference mapping. The spectral radius of $G$ is defined to be
@@
We recall that there exist simple numerical schemes able to compute the spectral radius [9, Remark 3][20]. In particular, if the general interference mapping is linear, then it can be written as ${\bm{x}}\mapsto{\bm{Mx}}$ for some nonnegative matrix ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$, and we can use classical results from Perron-Frobenius theory to confirm that $\rho({\bm{M}})$ in Definition 3, with $\bm{M}$} interpreted as a linear mapping, coincides with the usual notion of the spectral radius of the matrix $\bm{M}$} in linear algebra.
Having established the key terminology and concepts, we can proceed with our main contributions.
@@
To motivate the results in this section, consider the following optimization problem:
Example 1.
Let $K\in{\mathbb{N}}$, $N\in{\mathbb{N}}$, $x_{\mathrm{max}}>0$, ${\bm{a}}_{1},\ldots,{\bm{a}}_{K}]={\bm{A}}\in{\mathbb{R}}_{+}^{N\times K}$, $b_{1},\ldots,b_{N}]^{t}={\bm{b}}\in{\mathbb{R}}_{++}^{N}$, ${\bm{c}}_{1},\ldots,{\bm{c}}_{N}]={\bm{C}}\in{\mathbb{R}}_{+}^{N\times N}$, and $\sigma_{1},\ldots,\sigma_{N}]^{t}={\bm{\sigma}}\in{\mathbb{R}}_{++}^{N}$ be given problem parameters. For notational simplicity, define ${\bm{M}}:=\mathrm{diag}({\bm{b}})^{-1}{\bm{C}}^{t}$, ${\bm{u}}:=\mathrm{diag}({\bm{b}})^{-1}{\bm{\sigma}}$, and $(\forall n\in\{1,\ldots,K\})~{\bm{M}}_{n}:={\bm{M}}+(1/x_{\mathrm{max}}){\bm{u}}{\bm{a}}_{n}^{t}$. Assume that the function ${\mathbb{R}}^{N}\to{\mathbb{R}}_{+}:{\bm{x}}\mapsto\dfrac{1}{x_{\mathrm{max}}}\max_{n\in\{1,\ldots,K\}}{\bm{a}}_{n}^{t}|{\bm{x}}|=:\|{\bm{x}}\|$ is a polyhedral monotone norm, and consider the optimization problem:
$\begin{array}[]{rl}\text{maximize}_{{\bm{x}}=(x_{1},\ldots,x_{N})\in{\mathbb{R}}^{N}}&\min_{k\in\{1,\ldots,N\}}\dfrac{b_{k}x_{k}}{{\bm{c}}_{k}^{t}{\bm{x}}+\sigma_{k}}\\ \text{s.t.}&(\forall n\in\{1,\ldots,K\})~{\bm{a}}_{n}^{t}{\bm{x}}\leq x_{\mathrm{max}}\\ &{\bm{x}}\geq{\bm{0}}.\end{array}$ (8)
Standard results in the literature [35, Ch. 5.6.3][26, Proposition 4] show that a (not necessarily unique) solution ${\bm{x}}^{\star}\in{\mathbb{R}}_{++}^{N}$ to Problem (8) is the eigenvector (normalized to $\|{\bm{x}}^{\star}\|=1$) associated with the largest eigenvalue of any matrix ${\bm{M}}_{n^{\star}}$ with index $n^{\star}\in\{1,\ldots,K\}$ satisfying $\rho({\bm{M}}_{n^{\star}})=1/t^{\star}$, where $t^{\star}:=1/(\max_{n\in\{1,\ldots,K\}}\rho({\bm{M}}_{n}))$ is the optimal objective of (8).
Readers familiar with wireless systems may recognize the fractions in the objective function of Problem (8) as SINR expressions for radio links [26][35, Ch. 5.6.3][23, Ch. 5.3.2][2, Theorem 7.1], although we postpone a detailed discussion of these connections (in more general settings) to the next sections. From a high-level perspective, the characterization of the solution to Problem (8) provided in Example 1 is somewhat cumbersome: it requires comparing the spectral radii of an entire family of matrices. As a first concrete application of the main results of this section, we show that the same solution can instead be expressed more naturally through the eigenpair a possibly nonlinear general interference mapping. In particular, we later establish the following alternative characterization:
Remark 1.
Define a (possibly nonlinear) general interference mapping by $T_{\|\cdot\|}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}:{\bm{x}}\mapsto{\bm{Mx}}+{\bm{u}}\|{\bm{x}}\|$, where $\bm{M}$}, $\bm{u}$}, $x_{\mathrm{max}}$, and $\|\cdot\|$ are as in Example 1. A solution to Problem (8) is the eigenvector ${\bm{x}}^{\star}\in{\mathbb{R}}_{++}^{N}$ of $T_{\|\cdot\|}$}, normalized to satisfy $\|{\bm{x}}^{\star}\|=1$, associated with the nonlinear spectral radius $\rho(T_{\|\cdot\|})>0$, and $1/\rho(T_{\|\cdot\|})$ is the optimal objective of Problem (8).
The above nonlinear viewpoint has several advantages. It not only allows us a compact description of solutions to many problems in the wireless literature, but it also enables us to extend the matrix-based results of Example 1 to genuinely nonlinear scenarios, such as those encountered in wireless systems employing optimal beamforming strategies [27]. Furthermore, it enables the use of existing numerical methods for computing nonlinear eigenvectors to solve variants of Problem (8), which may prove valuable in future research. Most importantly for the developments that follow, the framework we develop provides the mathematical machinery necessary to analyze the geometry of utility regions in wireless networks.
To derive the results mentioned in Remark 1, we begin by recalling a standard fact from the wireless literature: the pointwise minimum of a finite family of standard interference functions results in a standard interference function [42]. However, this closure property does not generally extend to the pointwise infimum of an infinite family. A simple counterexample is given by the constant functions $f_{c}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}:{\bm{x}}\mapsto c$ parametrized by $c>0$. Each function $f_{c}$ with $c>0$ is a standard interference function, but the pointwise infimum $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})~f({\bm{x}}):=\inf_{c>0}f_{c}({\bm{x}})=0$ fails to satisfy the scalability property. Nevertheless, in wireless systems, the families of functions typically possess additional structure – such as that described in the next lemma – ensuring that the pointwise infimum is again a standard interference function.
@@
Let $\mathcal{Y}$ be a nonempty set. Suppose that, for each $y\in\mathcal{Y}$, $h_{y}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$ is a general interference function and ${u}_{y}\in{\mathbb{R}}_{++}$ is a positive scalar. Assume that there exists $\delta>0$ satisfying $(\forall y\in\mathcal{Y})~{u}_{y}\geq{\delta}>{0}$; i.e., the set $\{u_{y}\in{\mathbb{R}}_{++}~|~y\in\mathcal{Y}\}\subset{\mathbb{R}}_{++}$ is bounded away from zero. Define a function $t:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}$ by
@@
Assume that $t$ and $t_{\|\cdot\|}$ are continuous on ${\mathbb{R}}_{+}^{N}$. Then $t$ in (9) is a standard interference function, and $t_{\|\cdot\|}$ in (10) is a general interference function.
@@
We use the functions in (9) and (10) in Lemma 1 to define the following proper subclass of standard and general interference mappings: